2 points It is known that the population standard deviation of students debts is 11800 NIS. What is the minimum sample size2 points It is known that the population standard deviation of students debts is 11800 NIS. What is the minimum sample size
Answer #1

Solution(1)
Given in the question
Population standard deviation(\sigma) = 11800
The margin of error (E)= 800
Confidence level = 0.95
level of significance = 0.05, alpha/2 = 0.025
From Z table we found Zalpha/2 = 1.96
So sample size n can be calculated as
Sample size n = (Zalpha/2 * \sigma /E)^2 = (1.96*11800/800)^2 = 835.7881 or 836
The minimum sample size should be 836 so that the estimate that 95% confidence is within 800 NIS.
Solution(2)
Quartile 3 can be calculated as
Q3 = 3*(n+1)/4
Here n = 3*(8+1)/4 = 6.75th value i.e. from the data set 6th value is 44 and 7th value is 50 so 6.75th value is 44+0.75*(50-44) = 48.5
So Q3 = 48.5